A 0.50L sealed flask contains only 0.050g of glucose and 0.20 atm of oxygen. The glucose is ignited and burned entirely to carbon dioxide and water via the balanced equation below. Once the flask is cooled back to 298K, what is the total pressure within the system? (assume all of the water vapour remains gas)


C6H12O6 + 6O2 --> 6CO2 + 6H2O


1
Expert's answer
2021-01-05T05:56:39-0500

Moles of oxygen in flask = 0.0041 (Using Gas Law)

Moles of glucose in flask = 0.0028 ( Using molar mass and mass given)

Therefore,

Oxygen is limiting reagent. (from stoichiometry comparison)

Amount of glucose used up = 0.0007

Now from stoichiometry of equation it is evident that,

no. of moles of oxygen used = no. of moles of carbon dioxide released

   = no. of moles of water vapour released


Using gas law,

"P*V = n*R*T"


Here, T = 298 K , V = 0.5L , R = 0.082 L atm K-1 mol-1, n= 2 * 0.0041

"\implies" P = 0.40 atm